# Geometric Embeddability of Complexes is $\exists\mathbb{R}$-complete

### 05 Aug 2021 v1

@@ 1.2 Definitions
A $k$-simplex is a $k$-dimensional polytope which is the convex hull of its $k+1$ vertices. Hence, a 0-simplex corresponds to a point, a 1-simplex to a segment, and a 2-simplex to a triangle etc. The convex hull of any nonempty proper subset of the $k+1$ points that define a $k$-simplex is called a face of the simplex. A simplicial complex $K$ is a set of simplices satifying the following two conditions: (i) Every face of a simplex from $K$ is also in $K$. (ii) For any two simplices $\sigma_{1},\sigma_{2}\in K$ with a non-empty intersection, the intersection $\sigma_{1}\cap\sigma_{2}$ is a face of $\sigma_{1}$ and $\sigma_{2}$. The purely combinatorial counterpart to a simplicial complex is an abstract simplicial complex, which we refer to simply as a complex.
@@
A complex $C=(V,E)$ is a finite set $V$ together with a collection of subsets $E\subseteq 2^{V}$ which is closed under taking subsets, i.e., $e\in E$ and $e^{\prime}\subseteq e$ imply that $e^{\prime}\in E$. A $k$-complex is a complex where the largest subset contains exactly $(k+1)$ elements. We call a complex pure if all (inclusion-wise) maximal elements in $E$ have the same cardinality.
The neighbors of every vertex $v\in V$ in a $k$-complex $C=(V,E)$ define a lower dimensional complex $C_{v}:=(V^{\prime},E^{\prime})$ where $E^{\prime}:=\{e\setminus\{v\}|v\in e\in E\}$ and $V^{\prime}:=N(v)=\bigcup_{e\in E^{\prime}}e$ are the neighbors of $v$.
Complexes are in close relation to Hypergraphs.
@@
Hypergraphs generalize graphs by allowing edges to contain any number of vertices. Formally, an undirected hypergraph $H$ is a pair $H=(X,E)$ where $X$ is a set of vertices, and $E$ is a set of non-empty subsets of $X$ called hyperedges (or edges). A $k$-uniform hypergraph is a hypergraph such that all its hyperedges contain exactly $k$ elements. Note that the maximal sets of a pure $k$-complex yield a $(k+1)$-uniform hypergraph and vice versa. Hence, $(k+1)$-uniform hypergraphs and pure $k$-complexes are in a straight-forward one-to-one correspondence. A simplicial representation of a $(k+1)$-uniform hypergraph is a geometric embedding of the corresponding complex.
@@
PL-embeddings and topological embeddings.
In contrast to geometric embeddings, for PL or topological embeddings it is not sufficient to describe the mapping of the vertices. Thus, we define the embedding of a complex by its corresponding simplicial complex.
A PL/topological embedding of a simplicial complex $C=(V,E)$ in $\mathbb{R}^{d}$ is a piecewise/continuous linear function $\varphi\colon E\rightarrow\mathbb{R}^{d}$ such that the following property is satisfied: for every pair $e,e^{\prime}\in E$, it holds that
$\varphi(e)\cap\varphi(e^{\prime})=\varphi(e\cap e^{\prime}).$
@@
A pseudoline arrangement is a family of curves that apart from the ‘straightness’ share similar properties with a line arrangement. More formally, we define a (Euclidean) pseudoline arrangement as an arrangement of $x$-monotone curves in the Euclidean plane such that every pair of pseudolines meets exactly once in a crossing point. In fact, each pseudoline arrangement can be nicely encoded by a wiring diagram; see also Figure 2(a). A pseudoline arrangement is stretchable if it is combinatorially equivalent to an arrangement of straight-lines, i.e., if both arrangements are isomorphic. Stretchability denotes the algorithmic problem of deciding whether a given pseudoline arrangement is stretchable. In a seminal paper, Shor [50] proved that Stretchability is complete for the existential theory of the reals; for a stream-line exposition see Matoušek [33].
@@ 2 The Proof
Our proof of Theorem 1 consists of three parts.
@@
    We establish $\exists\mathbb{R}$-membership. (Lemma 3, Section 2.1)
@@
    We show $\exists\mathbb{R}$-hardness in $\mathbb{R}^{3}$, i.e., of $\textsc{GEM}_{2\rightarrow 3}$ and $\textsc{GEM}_{3\rightarrow 3}$. (Lemmas 4 and 5, Section 2.2)
@@
    We reduce $\textsc{GEM}_{k\rightarrow d}$ to $\textsc{GEM}_{k+1\rightarrow d+1}$. (Lemma 6, Section 2.3)
@@ 2.1 Membership
In this subsection, we show $\exists\mathbb{R}$-membership of $\textsc{GEM}_{k\rightarrow d}$. Note that this is essentially folklore [13]. We give here a proof for the sake of completeness.
@@
In order to show membership in $\exists\mathbb{R}$, we use the following characterization by Erickson, Hoog and Miltzow [22]: A problem $P$ lies in $\exists\mathbb{R}$ if and only if there exists a real verification algorithm $A$ for $P$ that runs in polynomial time on the real RAM. In particular, for every yes-instance $I$ of $P$ there exists a polynomial sized witness $w$ such that $A(I,w)$ returns yes, and for every no-instance $I$ of $P$ and any witness $w$, $A(I,w)$ returns no. In contrast to the definition of the complexity class NP, we also allow witnesses that consist of real numbers. Consequently, we execute $A$ on the real RAM as well.
It remains to present a real verification algorithm for $\textsc{GEM}_{k\rightarrow d}$. While the witness describes the coordinates of the vertices, the algorithm checks for intersections between any two simplices. Note that each simplex is a convex set and the intersection is a convex set as well. If the simplex $S$ has $n$ vertices, then we can efficiently determine $n$ linear inequalities and at most one linear equality that together describe $S$. Now checking for intersections can be reduced to linear programming, which is polynomial time solvable. This finishes the description of the real verification algorithm. ∎
@@ 2.2 Hardness in three dimensions
This section is dedicated to proving Theorem 1 for $k=2,3$ and $d=3$. The crucial part lies in the case $k=2$.
Lemma 4.
@@
We reduce from the $\exists\mathbb{R}$-hard problem Stretchability, as described in Section 1.2. Let $L$ be an arrangement of $n$ pseudolines in the plane. Every pseudoline arrangement has a representation as a wiring diagram where each pseudoline is represented by a monotone curve consisting of 2n12n-1 segments; for an illustration consider Figure 2(a). We add a pseudoline at infinity $\ell_{0}$ that intersects all pseudolines in the beginning. We call the resulting pseudoline arrangement $L^{*}$. Note that $L^{*}$ is stretchable if and only if $L$ is stretchable. In the following, we construct a 2-complex $C=(V,E)$ that allows for a geometric realization if and only if $L^{*}$ (and thus $L$) is stretchable.
@@
In order to define $C$, we use the intended geometric embedding. We will refer to the subsets in $C$ as vertices, edges, and triangles if they contain one, two or three elements, respectively. The construction has three steps.
Firstly, we place the pseudolines in 3-space such that each pseudoline $\ell_{i}$ lies the plane $z=i$. In particular, note that no two pseudolines intersect. Therefore, we can surround each (lifted) pseudoline by a triangulated tube; see also Figure 2(b).
(a) A pseudoline arrangement $L^{*}$.
(b) Construction of a tube.
(c) A section of a tube.
[…]
The tube is formed by $2n+1$ sections such that every second section covers a crossing (when viewed from the top). Moreover, each section consists of six triangles forming a triangulated triangular prism as illustrated in Figure 2(c). We close the tube with triangles at the ends and think of the bottom side of the prism to lie in the plane $z=i$. Hence, the tube has a distinguished top polyline.
Secondly, for pseudoline $\ell_{i}$, we insert a special edge $e_{i}$ between the two top vertices on either end of the tube; for later reference, we denote the end vertices by $s_{i}$ and $t_{i}$. The special edge $e_{i}$ is intended to lie inside the tube and represent the stretched pseudoline (if possible). To this end, we add a new vertex to the construction that we call apex and denote by $u$. Specifically, the edges of the triangles incident to $s_{i}$ and $t_{i}$ form a triangle with $u$; for an illustration, consider Figure 3.
(a) front view
(b) top view
(c) tent of $s_{i}$
Figure 3: Construction of the triangles incident to $u$ and edges ’close’ to $s_{i}$ (or $t_{i}$); orange edges form triangles with $u$.
Thirdly, triangles with the apex $u$ and almost every edge of the tubes will ensure the correct combinatorics of $L$. The idea is that the apex forms a triangle with every edge that it sees from the top (at infinity). Note that the apex does not see the bottom side of any tube and that an invisible top part of a tube corresponds to a crossing in $L$. Figure 4(a) illustrates the neighborhood of the apex at crossings.
(a) apex triangles at crossings
(b) Neighborhood complex $\overline{C}_{u}$ corresponds to a 3-connected planar graph.
Figure 4: Neighborhood of the apex (around crossings)
As of now, the neighborhood of the apex $u$ induces a plane 2-connected graph $C^{\prime}_{u}$. Lastly, we add one vertex in each face of $C^{\prime}_{u}$ which is not a triangle, see Figure 4(b); note that these faces are in bijection with the faces of $L^{*}$. We call the now constructed complex $\overline{C}$. Consequently, the neighborhood $N(u)$ induces a $1$-complex $\overline{C}_{u}$ that corresponds to a $3$-connected planar graph. Our final complex $C$ consist of two copies of $\overline{C}$ in which the apex vertices are identified. (We will later use these two copies in order to guarantee that the apex lies in the unbounded region for one copy of $\overline{C}$.)
@@
In order to verify that the construction shows $\exists\mathbb{R}$-hardness, we argue that it has a polynomial running time in the size of the input. To this end, note that a pseudoline arrangement with $n$ pseudolines can be described by the $O(n^{2})$ crossings. Thus,the input size is $N=O(n^{2})$. It is easy to see that our construction has a size linear in $N$: For each segment and crossing of the pseudoline arrangement, we build a constant size construction involving the apex. Consequently, the total construction has size $O(n^{2})=O(N)$.
@@
It remains to show that $L$ is stretchable if and only if $C$ has a geometric embedding in $\mathbb{R}^{3}$.
If $L$ is stretchable, the geometric embedding of $C$ is straight-forward and goes along the same lines as its construction: Considering the stretched line arrangement $L^{\prime}$ equivalent to $L$, we add $\ell_{0}$ and lift the lines in 3-space, construct the tunnels, make a second copy, and can insert the apex $u$ (high enough) above. This yields a geometric embedding of $C$.

### 05 Nov 2021 v2

@@ 1.2 Definitions
A $k$-simplex $\sigma$ is a $k$-dimensional polytope which is the convex hull of its $k+1$ vertices $V$, which are not contained in the same (k1)(k-1)-dimensional hyperplane. Hence, a 0-simplex corresponds to a point, a 1-simplex to a segment, and a 2-simplex to a triangle etc. The convex hull of any nonempty proper subset of $V$ is called a face of $\sigma$. A simplicial complex $K$ is a set of simplices satifying the following two conditions: (i) Every face of a simplex from $K$ is also in $K$. (ii) For any two simplices $\sigma_{1},\sigma_{2}\in K$ with a non-empty intersection, the intersection $\sigma_{1}\cap\sigma_{2}$ is a face of both simplices $\sigma_{1}$ and $\sigma_{2}$. The purely combinatorial counterpart to a simplicial complex is an abstract simplicial complex, which we refer to simply as a complex.
@@
A complex $C=(V,E)$ is a finite set $V$ together with a collection of subsets $E\subseteq 2^{V}$ which is closed under taking subsets, i.e., $e\in E$ and $e^{\prime}\subseteq e$ imply that $e^{\prime}\in E$. A $k$-complex is a complex where the largest subset contains exactly $k+1$ elements. We call a complex pure if all (inclusion-wise) maximal elements in $E$ have the same cardinality.
For any vertex $v\in V$ in a $k$-complex $C=(V,E)$, the neighbourhood of $v$ gives rise to a lower dimensional complex $C_{v}:=(V^{\prime},E^{\prime})$, where $E^{\prime}:=\{e\setminus\{v\}\mid v\in e\in E\}$ and $V^{\prime}:=N(v)=\bigcup_{e\in E^{\prime}}e$ are the neighbors of $v$. Complexes are in close relation to Hypergraphs.
@@
Hypergraphs generalize graphs by allowing edges to contain any number of vertices. Formally, a hypergraph $H$ is a pair $H=(V,E)$ where $V$ is a set of vertices, and $E$ is a set of non-empty subsets of $V$ called hyperedges (or edges). A $k$-uniform hypergraph is a hypergraph such that all its hyperedges contain exactly $k$ elements. Note that the maximal sets of a pure $k$-complex yield a $(k+1)$-uniform hypergraph and vice versa. Hence, $(k+1)$-uniform hypergraphs and pure $k$-complexes are in a straight-forward one-to-one correspondence. A simplicial representation of a $(k+1)$-uniform hypergraph is a geometric embedding of the corresponding complex.
@@
Topological and PL embeddings.
Consider a complex $C=(V,E)$. In contrast to geometric embeddings, for PL or topological embeddings it is not sufficient to describe the mapping of the vertices $V$. Choose $d^{\prime}$ so large that $C$ admits a geometric embedding $\varphi^{\prime}:V\to\mathbb{R}^{d^{\prime}}$, and define $S=\bigcup_{e\in E}\overline{\varphi^{\prime}}(e)$. We then say that an injective and continuous function $\varphi:S\to\mathbb{R}^{d}$ is a topological embedding of $C$ in $\mathbb{R}^{d}$. If furthermore for each $e\in E$, the image $\varphi(\overline{\varphi^{\prime}}(e))$ is a finite union of connected subsets of (|e|1)(|e|-1)-dimensional hyperplanes, then $\varphi$ is a piecewise linear (PL) embedding. The problem $\textsc{Embed}_{k\rightarrow d}$ asks whether a given $k$-complex has a PL embedding in $\mathbb{R}^{d}$.
Graph Drawings.
A graph is a 1-complex. A graph is planar if there exists a crossing-free drawing in the plane, i.e., a (topological) embedding in $\mathbb{R}^{2}$. As mentioned above, a graph has a topological embedding in $\mathbb{R}^{2}$ if and only if it has a geometric embedding in $\mathbb{R}^{2}$. A plane graph is a planar graph together with a specified crossing-free drawing. By means of stereographic projection, any graph that has a crossing-free drawing in the plane also has a crossing-free drawing on the sphere and vice versa. Two drawings of a graph (in the plane or on the sphere) are equivalent if they can be transformed into one another by a homeomorphism (of the plane or the sphere). In particular, two equivalent drawings have the same set of faces. Consider a plane graph $G$ and let $D^{\prime}$ be the specified drawing of $G$. When talking about an (arbitrary) drawing $D$ of $G$, we always mean that $D$ is equivalent to $D^{\prime}$.
@@
A pseudoline arrangement is a family of curves that apart from ‘straightness’ share similar properties with a line arrangement. More formally, a (Euclidean) pseudoline arrangement is an arrangement of $x$-monotone curves in the Euclidean plane such that any two meet in exactly one point. In fact, each pseudoline arrangement can be encoded by a wiring diagram; see also Figure 4(a). A pseudoline arrangement is stretchable if it is combinatorially equivalent to an arrangement of straight-lines, i.e., if the arrangements can be transformed into one another by a homeomorphism of the plane. Stretchability denotes the algorithmic problem of deciding whether a given pseudoline arrangement is stretchable. In a seminal paper, Shor [undefaaa] proved that Stretchability is complete for the existential theory of the reals; for a stream-line exposition of this result see the expository paper by Matoušek [undefah].
1.3 Pitfalls
While the general proof ideas are fairly straightforward, our arguments in Section 2 may at first glance appear a bit tedious. In the following, we highlight one of the appearing challenges. It is easy to see that each special edge lies inside its tunnel in any geometric embedding. It follows that the projection of the special edge lies also inside the projection of the tunnel on the sphere centered at the apex. Furthermore, we know that the roof of the tunnels are seen by the apex. One may be tempted to (directly) conclude that the projection of the special edge is thus also contained in the projection of the roof; the underlying thought being that the projection of the tunnel bottom lies below the tunnel roof in the geometric representation and thus the projection of the tunnel bottom is contained in the projection of the tunnel roof. Yet, the latter is not true in general, as can be seen in Figure 3. In the figure, the tunnel bottom is not covered by the roof. We (implicitly) show that the projection of the special edge lies inside the projection of the roof by establishing some even stronger topological and geometric properties.
Figure 3: From the perspective of $u$, the tunnel bottom is not always hidden below the tunnel roof. From the three tunnel sections displayed, the bottom (yellow) of the middle section is partially visible from the apex.
@@ 2 The Proof
In this section, we prove Theorem 1. Our proof consists of the following three parts.
@@
    Establishing $\exists\mathbb{R}$-membership (Section 2.1: Lemma 3).
@@
    Showing $\exists\mathbb{R}$-hardness in $\mathbb{R}^{3}$, i.e., of $\textsc{GEM}_{2\rightarrow 3}$ and $\textsc{GEM}_{3\rightarrow 3}$ (Section 2.2: Theorems 4 and 5).
@@
    Reducing $\textsc{GEM}_{k\rightarrow d}$ to $\textsc{GEM}_{k+1\rightarrow d+1}$ (Section 2.3: Lemma 6).
@@ 2.1 Membership
In this subsection, we show $\exists\mathbb{R}$-membership of $\textsc{GEM}_{k\rightarrow d}$. Note that this is essentially folklore [undefm]. We present a proof for the sake of completeness.
@@
In order to show membership in $\exists\mathbb{R}$, we use the following characterization by Erickson, Hoog and Miltzow [undefv]: A problem $P$ lies in $\exists\mathbb{R}$ if and only if there exists a real verification algorithm $A$ for $P$ that runs in polynomial time on the real RAM. In particular, for every yes-instance $I$ of $P$ there exists a polynomial sized witness $w$ such that $A(I,w)$ returns yes, and for every no-instance $I$ of $P$ and any witness $w$, $A(I,w)$ returns no. In contrast to the definition of the complexity class NP, we also allow witnesses that consist of real numbers. Consequently, we execute $A$ on the real RAM as well.
It remains to present a real verification algorithm for $\textsc{GEM}_{k\rightarrow d}$. While the witness describes the coordinates of the vertices, the algorithm checks for intersections between any two simplices. Note that each simplex is a convex set and the intersection of convex sets is a convex set as well. For any simplex $S$ with $n$ vertices, we can efficiently determine $n$ linear inequalities and at most one linear equality that together describe $S$. Then checking for intersections can be reduced to a linear program, which is polynomial time solvable. This finishes the description of the real verification algorithm. ∎
@@ 2.2 Hardness in three dimensions
This section is dedicated to proving Theorem 1 for $d=3$ and $k\in\{2,3\}$. The crucial part lies in the case $k=2$. For the benefit of the reader, we include a glossary in Table 3.
Theorem 4.
@@
We reduce from the $\exists\mathbb{R}$-hard problem Stretchability, as described in Section 1.2. Let $L$ be an arrangement of $n$ pseudolines in the plane. Every pseudoline arrangement has a representation as a wiring diagram in which each pseudoline is given by a monotone curve consisting of 2n12n-1 sections. For an illustration consider Figure 4(a); each section could be represented by a segment, however for a visual appealing display, the bend points are rounded. We add a pseudoline $\ell_{0}$ that intersects all pseudolines in the beginning, see Figure 4(a), and call the resulting pseudoline arrangement $L^{*}$. Note that $L^{*}$ is stretchable if and only if $L$ is stretchable. For later reference, we endow a natural orientation upon each pseudoline from left to right.
(a) A pseudoline arrangement $L^{*}$.
(b) The crossing diagram contains an additional helper triangle.
Figure 4: Adding an extra pseudoline $\ell_{0}$ and the helper triangle $\triangle$ to the construction.
In the following, we construct a 2-complex $C=(V,E)$ that allows for a geometric realization if and only if $L^{*}$ (and thus $L$) is stretchable. In order to define $C$, we add a helper triangle $\triangle$ to our arrangement that intersects the pseudolines of $L^{+}$ as illustrated in Figure 4(b).
@@
In order to define $C$, we use the intended geometric embedding. We will refer to the subsets in $C$ as vertices, edges, and triangles depending on whether they contain one, two or three elements. The construction has five steps.
In the first step, we place the pseudolines and the helper triangle $\triangle$ in 3-space. Each pseudoline $\ell_{i}$ lies in the plane $z=i$ such that an observer high above (at infinity) sees the wiring diagram. Similarly, we place the segments of the helper triangle $\triangle$ in 3-space such that it lies in the plane $z=n+1$. Note that no two pseudolines intersect. Therefore, we can surround each lifted pseudoline by a triangulated sphere which we call a tunnel; see also Figure 5(a). The tunnel $T^{+}_{i}$ of $\ell_{i}$ is formed by $2n+3+i$ sections; later, we will be particularly interested in a part of a tunnel, denoted by $T_{i}$, in which the first two and last two sections are removed. Each section consists of six triangles forming a triangulated triangular prism as illustrated in Figure 5(b). We close the tunnel with triangles at the ends and think of the bottom side of the prism to lie in the plane $z=i-\nicefrac{{1}}{{2}}$ (for now). The remaining part of the tunnel, i.e., the tunnel without its bottom, constitutes the roof, see Figure 5(c). The roof contains three disjoint paths of length $2n+4+i$. The edges and vertices on the boundary of the bottom and the roof form the left and right roof path, respectively, when deleting the edges of the closing triangles. The remaining vertices induce the central roof path. The three roof paths are thickened in Figures 5(a) and 5(c).
(a) A tunnel viewed from side.
(b) A tunnel section.
(c) A tunnel viewed from above.
Figure 5: First step in the construction of the complex $C$ – tunnel construction.
Note that we do not add a tunnel for the helper triangle. We distribute the sections along $T^{+}_{i}$ to edges and crossings of the crossing diagram as follows: Generally, we associate one section per edge and one section per crossing of two pseudolines. Moreover, we associate one extra section of $T^{+}_{j}$ to a crossing of $\ell_{i}$ and $\ell_{j}$ whenever $i<j$. In order to represent the pseudoline $\ell_{i}$, we insert a special edge $e_{i}$ between the two top vertices on either end of the tunnel; for later reference, we denote the start vertex by $s_{i}$ and the end vertex by $t_{i}$. The special edge $e_{i}$ is intended to lie inside the tunnel.
In the second step, we identify parts of the tunnels. To this end, consider the tunnel sections assigned to a crossing of a pseudoline $\ell_{i}$ with $\ell_{j}$, $i<j$. Recall that we assigned one section of $T^{+}_{i}$ and two sections of $T^{+}_{j}$ to the crossing. We identify the four triangles in the bottom of the two sections of $T^{+}_{j}$ with the four triangles in the roof of one section of $T^{+}_{i}$ as indicated in Figure 6. Note that we hereby identify six vertices, four of which belong to a left or right roof path of both, $T^{+}_{i}$ and $T^{+}_{j}$.
Figure 6: Second step in the construction of the complex $C$: gluing of tunnel parts.
In the third step, we add a new vertex to the construction that we call the apex and denote it by $u$. We think of $u$ as the observer high above (at infinity) and insert a triangle defined by $u$ and the vertices of every tunnel edge that is visible from $u$. Note that every roof section that is neither glued in a crossings nor hidden by the helper triangle is visible. Moreover, no bottom of any tunnel is visible in the intended geometric embedding.
In the fourth step, we enhance the $1$-complex induced by the neighborhood $N(u)$ of the apex $u$ such that it corresponds to an essentially $3$-connected planar graph $G^{+}$. We call a graph essentially $3$-connected if it is a subdivision of a $3$-connected graph. With the description so far, the $1$-complex corresponds to the graph $H$ depicted in black in Figure 7.
Figure 7: Third and fourth step in the construction of the complex $C$: neighborhood of the apex $u$. The graph $H$ depicted in black after the third step. Together with the gray edges, the graph is a candidate for the essentially 3-connected plane graph $G^{+}$ and it subgraph $G$ inside $\triangle$.
To construct $G^{+}$, we make use of the following fact.
Claim 1.
For every plane graph $G_{1}=(V_{1},E_{1})$, there exists an essentially 3-connected plane graph $G_{2}=(V_{2},E_{2})$ such that $G_{1}$ is a subgraph of $G_{2}$ and any straight-line drawing $D_{1}$ of $G_{1}$ in the plane can be extended to a straight-line drawing of $G_{2}$. Moreover, if the maximum face degree of $G_{1}$ is $k$, then the size of $G_{2}$ can be bounded by $|V_{2}|+|E_{2}|\leq O(k|V_{1}|)$.
Proof.
In order to construct $G_{2}$, we start with a drawing of $G_{1}$. First, we ensure that $G_{1}$ has at least 4 vertices. Second, we guarantee 2-connectedness by inserting edges; the edges are represented by potentially non-straight curves. To this end, we iteratively insert edges between different connected components; this ensures connectedness. Afterwards we insert one vertex in each face and insert one edge to each incident vertex of this face; this ensures 2-connectedness. Third, we triangulate each face by repeating the last step: we insert a vertex into each face and insert an edge to every vertex incident to this face. By 2-connectedness, we obtain a triangulation $T$ on at least four vertices and hence, a 3-connected plane graph. Lastly, we subdivide each new edge so that the number of subdivision vertices equals to the degree of the face in $G_{1}$ in which the edge has been inserted. The resulting graph is $G_{2}$ and by construction essentially 3-connected. Note that $T$ has $O(|V_{1}|)$ vertices. Therefore, $G_{2}$ has $O(k|V_{1}|)$ edges.
It remains to show that any straight-line drawing $D_{1}$ of the plane graph $G_{1}$ can be extended to a straight-line drawing of the planar graph $G_{2}$. Because $D_{1}$ and $G_{1}$ have the same set of faces, we can insert the additional edges of $T$ by polylines. Following the face boundary, the number of bends on each edge is upper bounded by the degree of the face of $G_{1}$ in which it is inserted. Hence, we can easily extend $D_{1}$ to a straight-line drawing of $G_{2}$. ∎
Let $G^{+}:=G_{2}$ be an essentially 3-connected plane graph guaranteed by Claim 1 for the case that $G_{1}=H$. Note that $G_{1}$ has $O(n^{2})$ vertices and edges, and every face has degree $O(n)$. Hence, the size of $G_{2}$ is in $O(n^{3})$. We denote the outer face of $G_{2}$ by $f_{0}$. The reader is invited to think about the far more sparse graph depicted in Figure 7, which also serves as a candidate for $G^{+}$. Indeed, the depicted graph also fulfills all properties necessary for our construction; however, not all properties of Claim 1. For example, the depicted graph is even $3$-connected. The proof of this is straightforward, but a bit tedious. Thus, we leave it as an exercise to the interested reader to check that the graph remains connected even after the deletion of any two vertices or alternatively, that any pair of vertices is connected by three disjoint paths.
Later, the subgraph $G$ of $G^{+}$ that is induced by all vertices of $\bigcup_{i}T_{i}$ will be of particular interest; in Figure 7, these vertices (and their convex hull) lie inside the helper triangle $\triangle$. Recall that $T_{i}$ denote the part of the tunnel $T^{+}_{i}$ obtained by deleting the first two and last two sections.
It is a well-known fact that all (straight-line or topological) planar drawings of a 3-connected planar graph on the sphere are equivalent [undefae]; for a definition of equivalent drawings consult Section 1.2. Consequently, the result extends to essentially $3$-connected graphs as it also holds for topological drawings. For later reference, we note the following.
Claim 2.
The planar graph $G^{+}$ is essentially $3$-connected. Therefore, all crossing-free drawings of $G^{+}$ on a sphere are equivalent. Furthermore, any straight-line drawing of $H$ in the plane can be extended to a straight-line drawing of $G^{+}$.
We ensure that the neighborhood complex of $u$ is the underlying planar graph of $G^{+}$, i.e, for each edge of $G^{+}$ not present in $H$, we insert a triangle formed by the vertices of this edge together with $u$ and call the resulting complex $\overline{C}$.
In the fifth and last step, our final complex $C$ consist of two copies of $\overline{C}$ in which the apex vertices are identified. We will later use these two copies in order to guarantee that in any geometric embedding the apex lies outside of all tunnels for one copy of $\overline{C}$. This finishes the construction of the complex $C$.
@@
In order to verify that the construction shows $\exists\mathbb{R}$-hardness, we argue that it has a running time that is polynomial in the size of the input. To this end, note that a pseudoline arrangement with $n$ pseudolines can be described by the order of the $O(n^{2})$ crossings. Thus, the input size is $N=O(n^{2})$. After adding the helper triangle and $\ell_{0}$, the crossing diagram still has a size in $O(n^{2})$. It is easy to see that our construction has a size proportional to $N^{3/2}$: For each segment and crossing of the diagram, we build a constant size construction involving the apex. Moreover, we add a triangle for every (additional) edge in $G^{+}$; recall that $G^{+}$ has size $O(n^{3})$. Consequently, the total construction has size $O(n^{3})=O(N^{3/2})$. We remark, that a more careful choice of $G^{+}$, as in Figure 7, yields a construction that is linear in $N$.
@@
It remains to show that the pseudoline arrangement $L$ is stretchable if and only if $C$ has a geometric embedding in $\mathbb{R}^{3}$.
Correctness I: Stretchability implies Embedabbility.
If $L$ is stretchable, it is relatively straight-foward to construct a geometric embedding of $C$.
Claim 3.
If $L$ is stretchable, then $C$ has a geometric embedding.
Proof.
The construction of the geometric embedding goes along the same lines as the construction of $C$. Consider the stretched line arrangement $L^{\prime}$ equivalent to the pseudoline arrangement $L$, we construct $L^{\prime\prime}$, by adding $\ell_{0}$ to the left of all crossings of $L^{\prime}$.
Next we show how to add the helper triangle $\triangle$ such that the intersection pattern is as depicted in the crossing diagram. Note that we merely have to construct the triangle in a way that all intersections of $L^{\prime\prime}$ are contained in it and the vertices of the helper triangle lie in the correct faces. For an illustration of the following argument see Figure 8.
Figure 8: Illustration for the proof of Claim 2 concerning the insertion of $\triangle$. By scaling $L^{\prime\prime}$ while keeping point $p$ and the helper triangle fixed, we can ensure that all intersections lie inside the helper triangle eventually.
The helper triangle has three vertices $v_{0},v_{1},v_{2}$. By construction, the two vertices $v_{0},v_{1}$ are supposed to lie in the bottom cell of $L^{*}$ which is bounded by $\ell_{0}$ and $\ell_{n}$ and potentially other pseudolines. Similarly, $v_{2}$ is in the upper cell bounded exactly by $\ell_{0}$ and $\ell_{n}$. Let us fix some triangle $\triangle$ that contains the intersection point $p$ of $\ell_{0}$ and $\ell_{n}$ with the properties above. Consider the (gray) region that contains all intersections of $L^{\prime\prime}$. By scaling $L^{\prime\prime}$ while fixing $p$ and the triangle $\triangle$, the gray region becomes arbitrarily small and is eventually contained in $\triangle$. This shows that the helper triangle can be added as wished.
Afterwards, we lift the lines and the helper triangle in $3$-space, construct the tunnels, and enhance $H$ to $G^{+}$. For the latter step, we use the fact that we can extend any straight-line drawing of $H$ to a straight-line drawing of $G^{+}$ by Claim 2. (In case that we want to use $G^{+}$ as in Figure 7, we need to use the fact that the drawing of $H$ comes from the stretched line arrangement and thus the faces are convex.) Afterwards, we insert the apex $u$ (high enough) above, make a second copy by taking the mirror image, and identify the apices. This yields a geometric embedding of $C$. ∎
Correctness II: Embedabbility implies Stretchability.
The reverse direction is more involved. Let $\varphi$ denote a geometric embedding of $C$. To show that $L$ is stretchable, we start with a collection of crucial properties.